OpenStudy (sleepyjess):

Okay, so I was looking at my College Algebra book for next semester, and I'm confused on at least one thing so far. Graphing a sequence, can anyone help?

OpenStudy (sleepyjess):

The example in the book is: Write down the first six terms of the following sequence and graph it. \(a_n = \dfrac{n-1}n\) How do I know what the x and y values are?

OpenStudy (sleepyjess):

So for a = 1 \(a_1 = \dfrac{1-1}1 = 0\) Would x be 1 and y be 0? Making the coordinate (1, 0)?

OpenStudy (sleepyjess):

Then \(a_2 = \dfrac{2-1}2 = \dfrac12\) x = 2 and y = 1/2? Making the coordinate (2, 1\2)?

OpenStudy (sleepyjess):

Do I have the right idea or nah? :)

OpenStudy (shadowlegendx):

You read your books before school starts? Jess you tryhard

OpenStudy (sleepyjess):

lol, I like math and I haven't taken it yet this year, so I was refreshing :P

OpenStudy (mathmale):

\[a _{n}=\frac{ x-1 }{ n }\]

OpenStudy (mathmale):

There are no x or y here, although you can graph this sequence. Just let n begin with the value 1 and increase it: 1, 2, 3, 4, ..... The first 2 terms are \[\frac{ 2-1 }{ 2 },\frac{ 3-1 }{ 3 }\]

OpenStudy (mathmale):

These terms can, of course, be simplified.

OpenStudy (sleepyjess):

What would I use for coordinates to graph it?

OpenStudy (sleepyjess):

Wouldn't the first term be \(\dfrac01\)?

jimthompson5910 (jim_thompson5910):

@sleepyjess you are correct each point generated by the sequence is of the form (n, f(n)) f(n) is the function rule for the sequence, in this case, f(n) = (n-1)/n

jimthompson5910 (jim_thompson5910):

n is a natural number (it is a positive whole number)

OpenStudy (sleepyjess):

Thanks both of you :)